# Extrinsic Evaluation of Cultural Competence in Large Language Models

### v2

@@ 5.3 Correlation in Outputs & Cultural Values
When HCD is used as the measure of difference in cultural values (Figure 2(a)), we find that median correlation across the board121212except QA for Gemma 2B Instruct is greater than 0. This implies a small but positive correlation between the text distribution and cultural values of countries as measured by HCD. However, when WVS data is used, we find a small and negative correlation between text distribution and cultural values as measured by WVS (Figure 2(b)). All the rank correlation values were statistically significant within a significance interval of 95% in a two-sided p-test.
@@ 6 Discussion
Together the findings for RQ2 (§ 5.2) and RQ3 (§ 5.3) suggest that intrinsic and extrinsic measures of cultural competence are not correlated. On the one hand, model outputs from our extrinsic setup feature culturally relevant words (§ 5.2). On the other hand, the text distributions are only weakly correlated with measures of cultural values widely used in intrinsic evaluations of cultural competence (§ 5.3). Thus, even if an LLM reflects the values of every country perfectly (as prior work measures by Hofstede’s Cultural Dimensions or World Values Survey), this ability may not be reflective of cultural competence in downstream tasks. [^13] Complementary facets of intrinsic and extrinsic evaluation have been observed in multiple settings. For example, there is limited correlation between intrinsic and extrinsic fairness metrics Gonen and Goldberg (2019); Goldfarb-Tarrant et al. (2021); Cao et al. (2022), and in intrinsic metrics of language model quality (like perplexity) and downstream task performanceFaruqui et al. (2016); Dudy and Bedrick (2020).
@@
We show that models adapt to explicit cues of culture by using culturally relevant words (§ 5.2). But, how this will affect user experience is unclear. In prior work, Lucy et al. (2023) found mixed reactions from users when an email auto-reply system adapted to cues of their identities. Moreover, we do not consider any implicit cues of culture, like dialect or topical differences in queries Kirk et al. (2024). Thus, understanding whether model adaptations triggered by implicit and explicit cues of culture are useful or desired by users remains open.
Further, as the qualitative evaluation shows, the output contains names that are typically associated with the ethnic majorities of the country. This is reflective of biases of the models, which can also lead to potentially offensive, and hurtful generations. While user-facing LLMs might have some, albeit imperfect, safeguards against generating outright toxic content, they might still generate stereotypical text for marginalized groups and cause representational harms Gadiraju et al. (2023).
@@
We find that the correlation between text similarity and cultural values is affected by the measure of the cultural values (§ 5.3). One of the reasons for this might be that measures of cultural values like HCD and WVS are imperfect and incomplete. This is because are ample disagreements on the very definition of culture Baldwin et al. (2006). In fact, Hofstede’s Cultural Dimension Theory has been widely criticized for its static nature and over-simplification of culture Signorini et al. (2009). Even so, evaluating cultural competence in LLMs heavily relies on these measures of culture, inheriting these flaws. Future work should consider diverse and complimentary measures of culture.
Further, like the Hofstede’s theory, most evaluations of cultural competence are also done using static benchmarks. However, the world is an evolving place where cultural norms and values are not static. They change and develop through complex interactions among societies. Future work should focus on incorporating evaluation methods like dynamic benchmarking Kiela et al. (2021) or dealing with disagreements Davani et al. (2022), among others to account for the evolving nature of culture.
Finally, in our work, we use nationality as a proxy for culture. Our choice was motivated by the availability of data for cultural values for countries and by similar operationalization in prior work. However, culture cannot be anchored by nationalities alone. Moreover, countries are not monoliths and comprise of many and diverse communities. Calls for inclusive evaluations of fairness in language technologies Bhatt et al. (2022) have led to important recent work on building fairness resources with participatory design Dev et al. (2023b, a). We believe that methods of evaluation of cultural competence should also similarly embrace participatory and intersectional design.
@@ 7 Limitations
We hope that future work can address these limitations to holisitcally evaluate LLMs for cultural competence in user-facing tasks.
@@ 8 Conclusion
In this work, we evaluated cultural competence in two tasks, story generation and open-ended question answering. Our data contributions include a hand-curated list of 345 diverse question-answering topics and 35 story generation topics. We also obtain model outputs for 6 models and 195 nationalities which we will make available for further analysis. Our methodological contributions include conceiving two quantitative and one qualitative analyses for evaluation of LLM outputs for cultural competence. Using these methods, we find that models do vary their outputs with varying nationalities (§ 5.1), outputs contain culturally relevant artefacts (§ 5.2), and model outputs weakly correlate with cultural values (§ 5.3). Our findings underscore the importance of comprehensive extrinsic evaluation of cultural competence.
@@
We do not study any sensitive content in this paper, but we note that the outputs of the models could have potentially sensitive and offensive content. Further, the cultural competence of LLMs (or lack of thereof) can lead to varying experiences for users from different demographic backgrounds. We discuss the importance of considering user agency and representational harms in this context in § 6.
@@ Acknowledgements
We thank Saujas Vaduguru, Lucy Li, Arjun Subramonian, Jessica Huynh, Alfredo Gomez, To Eun Kim, Athiya Devyani, Simran Khanuja, Akhila Yerukola, Maarten Sap, Yulia Tsvetkov, and Sunipa Dev for their valuable feedback on the work. Special thanks to Daniel Voster, Dan Griffin, and Graham Neubig for their help in using the compute infrastructure.
Appendix A Datasheet
A.1 Motivation
This dataset has two parts. First is a list of topics to prompt models with for two tasks, question answering and story generation to analyse difference in model outputs across nationalities. Second are the model responses for these prompts.
Anonymized for peer review
What support was needed to make this dataset? (e.g.who funded the creation of the dataset? If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
Anonymized for peer review
A.2 Composition
The data consists of a list of topics. The model outputs contain text generated by a LLMs.
How many instances are there in total (of each type, if appropriate)?
35 topics for story generation and 345 topics for QA. For model outputs, each topic leads to 195 prompts (for 195 nationalities) and 5 responses are sampled for every prompt from 6 LLMs. This leads to 2018250 model outputs for QA and 175500 model outputs for stories.
This is a hand curated list of data. It is not exhaustively representative of all possible story generation topics or QA topics. For story generation in particular, we only focus on children’s stories. For QA, we attempt to include diverse topics and categories. But we note that these are open-ended tasks and thus the range of topics is very wide to measure exhaustiveness.
Each instance in the topic list is simply a phrase (unigram or bigram) that is used to create a prompt for Question answering or story generation. Each instance of model output is a paragraph with maximum 100 tokens in case of QA and 1000 tokens in case of story generation.
There are no labels
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
No
No
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
No
It is self-contained.
No
No
No
For collecting model outputs, the prompt that we use explicitly mention a nationality. This is because we want to study the perturbation of the model outputs when nationalities are perturbed in the prompts. Because of this model outputs in the data are likely to contain text that refer to respective nationalities.
No
No
Any other comments?
No
A.3 Collection
The topics were obtained by hand-curation. The authors first created a broad list of 13 categories that were of interest in the evaluation: biology, chemistry, environment, economics, history, humanities, law, maths, physics, politics, space, religion, world affairs. This categories were selected as intuitive categories of questions in which differences in model outputs might be observed. The authors then referred to textbooks and encyclopedia index to sample topics within these categories leading to a total 345 topics. For stories, the authors first similarly selected three broad categories on which children’s stories can be written: moral values, stories with specific characters, stories with specific settings. They then used online websites and children’s story books to come up with topics with these areas creating a list of 35 topics. This is the topic lists. Next, these were then used in a simple template ‘Explain {topic} to a / an {nationality} person.’ for QA and ‘Write a story about {topic} for a / an {nationality} kid.’ in story. The resulting prompts were input into 6 LLMs listed in 4.2 to obtain model outputs. 5 responses were generated for every output.
The entire data of topic list is human curated. The model outputs are LLMs generated. Some characteristics of the model outputs are evaluated in the paper.
The cost of hand-curating topic lists was about 10 researcher hours. For getting model outputs, A6000 GPUs was used for hosting the LLM to run inference for obtaining model outputs. The total inference cost was about 45 GPU hours. Model outputs from GPT 3.5 cost about 125 USD.
We did not sample.
The data was hand curated by the author and the author queried LLMs for model outputs. No additional personnel was involved.
No
NA
NA
NA
NA
NA
Any other comments?
NA
A.4 Preprocessing / Labelling / Cleaning
No
No cleaning was performed
NA
Any other comments?
NA
A.5 Uses
Yes. The data, paper, and code will be open sourced after peer review.
What (other) tasks could the dataset be used for?
No. We do note though that the model outputs are generated content from LLMs and might content toxic, offensive, and stereotypical texts against marginalized communties. We advise discretion on part of users who choose to further utlize this data for analysis.
The topic lists should not be treated an exhaustive list of topics to evaluate cultural competence. The model outputs should not be used as gold standard answers for the particular questions or story generation tasks.
Any other comments?
No
A.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
The data will be open-sourced
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
The data will be open-sourced onto a github repo or huggingface after publication.
When will the dataset be distributed?
The data will be open-sourced after publication.
The data will be open-sourced.
No
No
Any other comments?
YOUR ANSWER HERE
A.7 Maintenance
Who is supporting/hosting/maintaining the dataset?
Anonymized for peer review.
Anonymized for peer review.
Is there an erratum? If so, please provide a link or other access point.
No. The authors can be contacted via email.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
This data is unlikely to be updated.
NA
We do not intend to have multiple version.
TBD
Any other comments?
NA
Appendix B Lexical Variance
B.1 Calculation Details
@@
Within this equation $x_{i}-x_{j}$ essentially represents distance between the two points, which we replace with lexical distance or the Word Edit Distance (WED). Thus, repurposing the above variance equation, lexical variance in outputs across nationalities for a concept can be calculated as:
@@
B.2 ANOVA results on within and across nationality lexical variance
@@
Appendix C Kendalls Tau Rank Correlation
C.1 Choice of Kendalls Tau variant
@@
We use the c variant in particular because before the ranking both the rank list have been generated by metrics that have different scales.
C.2 An example calculation
@@
This is a brief example of how Kendall’s $\tau_{c}$ was calculated. Suppose there are 4 nationalities: A, B, C, D. We first take one nationality as an anchor, let’s say A, and create two rank lists. The first rank list is of similarity of text outputs to A, let’s say this is [B, D, C] and the second is using distance between cultural values representation (we reverse the raw rank list we get from distance in vector representation of cultural values, because this is distance while the other one similarity), let’s say this is [D, C, B]. For A, the rank correlation between these two ranklist is calculated using Kendall’s $\tau_{c}$ We use sklearn to calculate Kendall’s $\tau_{c}$ with default parameters. Finally, for a particular concept, we take average of Kendall’s $\tau_{c}$ across all nationalities.

### v3

@@ 5.3 Correlation in Outputs & Cultural Values
When HCD is used as the measure of difference in cultural values (Figure 6(a)), we find that median correlation across the board151515except QA for Gemma 2B Instruct is greater than 0. This implies a small but positive correlation between the text distribution and cultural values of countries as measured by HCD. However, when WVS data is used, we find a small and negative correlation between text distribution and cultural values as measured by WVS (Figure 6(b)). All the rank correlation values were statistically significant within a significance interval of 95% in a two-sided p-test.
@@ 6 Discussion
Together the findings for RQ2 (§ 5.2) and RQ3 (§ 5.3) suggest that intrinsic and extrinsic measures of cultural competence are not correlated. On the one hand, model outputs from our extrinsic setup feature culturally relevant words (§ 5.2). On the other hand, the text distributions are only weakly correlated with measures of cultural values widely used in intrinsic evaluations of cultural competence (§ 5.3). Thus, even if an LLM reflects the values of every country perfectly (as prior work measures by Hofstede’s Cultural Dimensions or World Values Survey), this ability may not be reflective of cultural competence in downstream tasks. [^16] Complementary facets of intrinsic and extrinsic evaluation have been observed in multiple settings. For example, there is limited correlation between intrinsic and extrinsic fairness metrics Gonen and Goldberg (2019); Goldfarb-Tarrant et al. (2021); Cao et al. (2022); Delobelle et al. (2022), and in intrinsic metrics of language model quality (like perplexity) and downstream task performance Faruqui et al. (2016); Dudy and Bedrick (2020).
@@
Our results show that models adapt to explicit cues of culture with culturally relevant words (§ 5.2). But, it is unclear how this will affect user experience. In prior work, Lucy et al. (2024) found mixed reactions from users when an email auto-reply system adapted to cues of their identities. Moreover, we do not consider any implicit cues of culture, like dialect or topical differences in queries Kirk et al. (2024). Understanding whether model adaptations triggered by implicit and explicit cues of culture are useful or desired by users remains open.
Further, as the qualitative evaluation shows, the output contains names that are typically associated with the ethnic majorities of the country. This is reflective of the biases of the models, which can also lead to potentially offensive, and hurtful generations. While user-facing LLMs might have some, albeit imperfect, safeguards against generating outright toxic content, they might still generate stereotypical text for marginalized groups and cause representational harms Gadiraju et al. (2023).
@@
We find that the correlation between text similarity and cultural values is affected by the measure of the cultural values (§ 5.3). One of the reasons for this might be that measures of cultural values like HCD and WVS are imperfect and incomplete. This is because there are ample disagreements on the very definition of culture Baldwin et al. (2006). In fact, Hofstede’s Cultural Dimension Theory has been widely criticized for its static nature and over-simplification of culture Signorini et al. (2009). Even so, evaluating cultural competence in LLMs heavily relies on these measures of culture, inheriting these flaws. Future work should consider diverse and complementary measures of culture.
Further, like Hofstede’s theory, most evaluations of cultural competence are also done using static benchmarks. However, the world is an evolving place where cultural norms and values are not static. They change and develop through complex interactions among societies. Future work should focus on incorporating evaluation methods like dynamic benchmarking Kiela et al. (2021) or dealing with disagreements Davani et al. (2022), among others to account for the evolving nature of culture.
Finally, in our work, we use nationality as a proxy for culture. Our choice was motivated by the availability of data on cultural values for countries and by similar operationalization in prior work. However, culture cannot be anchored by nationalities alone. Moreover, countries are not monoliths and comprise of many and diverse communities. Calls for inclusive evaluations of fairness in language technologies Bhatt et al. (2022) have led to important recent work on building fairness resources with participatory design Dev et al. (2023, 2024). We believe that methods of evaluation of cultural competence should also similarly embrace participatory and intersectional design.
@@ 7 Limitations
We hope that future work can address these limitations to holistically evaluate LLMs for cultural competence in user-facing tasks.
@@ 8 Conclusion
In this work, we evaluated cultural competence in two tasks, story generation and open-ended question answering. Our data contributions include a hand-curated list of 345 diverse question-answering topics and 35 story-generation topics. We also obtain model outputs for 6 models and 193 nationalities which we will make available for further analysis. Our methodological contributions include conceiving two quantitative and one qualitative analysis for the evaluation of LLM outputs for cultural competence. Using these methods, we find that models do vary their outputs with varying nationalities (§ 5.1), outputs contain culturally relevant artefacts (§ 5.2), and model outputs weakly correlate with cultural values (§ 5.3). Our findings underscore the importance of comprehensive extrinsic evaluation of cultural competence.
@@
We do not study any sensitive content in this paper, but we note that the outputs of the models could have potentially sensitive and offensive content. Further, the cultural competence of LLMs (or lack thereof) can lead to varying experiences for users from different demographic backgrounds. We discuss the importance of considering user agency and representational harms in this context in § 6. We note that the examples used in the paper are informed by the knowledge and lived experiences of the authors.
@@ Acknowledgements
We thank Saujas Vaduguru, Lucy Li, Arjun Subramonian, Anjali Kantharuban, Jessica Huynh, Alfredo Gomez, To Eun Kim, Athiya Devyani, Simran Khanuja, Akhila Yerukola, Jeremiah Milbauer, Maarten Sap, Yulia Tsvetkov, and Sunipa Dev for their valuable feedback throughout the progress of the work. Special thanks to Daniel Vosler, Dan Griffin, and Graham Neubig for their help and administration of the cluster compute.
Appendix A Lexical Variance
A.1 Calculation Details
@@
Within this equation, $x_{i}-x_{j}$ essentially represents the distance between the two points, which we replace with lexical distance or the Word Edit Distance (WED). Thus, repurposing the above variance equation, lexical variance in outputs across nationalities for a concept can be calculated as:
@@
A.2 ANOVA results on within and across nationality lexical variance
@@
(a) Story Generation
(b) Question Answering
Figure 5: Lexical Variance in outputs with temperature = 0.7. The variance of outputs across nationalities is consistently higher than the variance of outputs within nationalities, as also observed with a temperature of 0.3 in §5.1. Story generation has a higher median variance than QA across models. Note that the absolute values of variances across the board are higher than those obtained for the temperature = 0.3, which is consistent with the expectation of variation in generation increasing with increasing temperature.
Appendix B Kendalls Tau Rank Correlation
B.1 Choice of Kendalls Tau variant
@@
We use the c variant in particular because before the ranking both the rank lists have been generated by metrics that have different scales.
B.2 An example calculation
@@
This is a brief example of how Kendall’s $\tau_{c}$ was calculated. Suppose there are 4 nationalities: A, B, C, D. We first take one nationality as an anchor, let’s say A, and create two rank lists. The first rank list is of similarity of text outputs to A, let’s say this is [B, D, C] and the second is using the distance between cultural values representation let’s say this is [D, C, B]. Here, we reverse the raw rank list we get from distance in the vector representation of cultural values because this was obtained using distance values while the other one was obtained using similarity values, For A, the rank correlation between these two rank lists (i.e. [B, D, C] and [D, C, B]) is calculated using Kendall’s $\tau_{c}$ We use sklearn to calculate Kendall’s $\tau_{c}$ with default parameters. Finally, for a particular concept, we take an average of Kendall’s $\tau_{c}$ across all nationalities.
(a) Correlation with Hofstede’s Cultural Dimensions (HCD)
(b) Correlation with World Values Survey (WVS)
Figure 6: Kendall’s $\tau_{c}$ rank correlation between text distribution and cultural closeness of countries. For both plots, text similarity is measured using BLEU. For HCD correlation statistic values are greater than 0, implying a small but positive correlation (6(a)). However, for WVS, most correlations are less than 0, indicating small and negative correlation (6(b)). There are no clear trends among different models or tasks. We note that these findings are consistent with the findings of §5.3.
Appendix C Results with Temperature = 0.7
In this section, we describe the results obtained with outputs when a temperature of 0.7 is used during generation. We note that the findings are consistent with the results described in §5, albeit with different absolute values.
Figure 5 shows the lexical variance for outputs generated with temperature = 0.7. Consistent with findings from §5.1, the median variance across nationalities is higher than the median variance within nationalities, variance in story outputs is consistently more than QA outputs, and the difference between quartiles is higher for QA than for stories. All of this is consistent with findings from §5.1. Note that all the variance values across the board are higher than the variance values obtained with a temperature of 0.3. This is expected as randomness in generations of LLMs is known to increase with increasing temperature.
Next, for correlation between the similarity of model outputs and cultural values among nationality, we also observe consistent findings at a higher temperature. Specifically, similar to findings in §5.3, all Kendall’s $\tau_{c}$ correlation values are close to 0, generally on the +ve side when HCD is used and -ve when WVS is used, and there is no significant difference based on the task.
Appendix D Datasheet
D.1 Motivation
This dataset has two parts. First is a list of topics to prompt models with for two tasks, question answering and story generation to analyse differences in model outputs across nationalities. Second are the model responses for these prompts.
Anonymized for peer review
What support was needed to make this dataset? (e.g. funded the creation of the dataset? If there is an associated grant, provide the name of the grantor and the grant name and number, or if it was supported by a company or government agency, give those details.)
Anonymized for peer review
D.2 Composition
The data consists of a list of topics. The model outputs contain text generated by LLMs.
How many instances are there in total (of each type, if appropriate)?
35 topics for story generation and 345 topics for QA. We combine this with 193 nationalities to result in 66,585 QA prompts and 6755 Story prompts. We sample 5 responses for every prompt from 6 LLMs and two temperature settings (except GPT temperature 0.7). This leads to 3,662,175 model outputs for QA and 371,525 model outputs for stories.
This is a hand-curated list of data. It is not exhaustively representative of all possible story generation topics or QA topics. For story generation in particular, we only focus on children’s stories. For QA, we attempt to include diverse topics and categories. But we note that these are open-ended tasks and thus the range of topics is very wide to measure exhaustiveness.
Each instance in the topic list is simply a phrase (unigram or bigram) that is used to create a prompt for Question answering or story generation. Each instance of model output is a paragraph with a maximum of 100 tokens in case of QA and 1000 tokens in case of story generation.
There are no labels
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information but might include, e.g., redacted text.
No
No
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the the rationale behind them.
No
It is self-contained.
No
No
No
For collecting model outputs, the prompt that we use explicitly mentions a nationality. This is because we want to study the perturbation of the model outputs when nationalities are perturbed in the prompts. Because of this model outputs in the data are likely to contain text that refers to respective nationalities.
No
No
Any other comments?
No
D.3 Collection
The topics were obtained by hand-curation. The authors first created a broad list of 13 categories that were of interest in the evaluation: biology, chemistry, environment, economics, history, humanities, law, maths, physics, politics, space, religion, and world affairs. These categories were selected as intuitive categories of questions in which differences in model outputs might be observed. The authors then referred to textbooks and encyclopedia indexes to sample topics within these categories leading to a total of 345 topics. For stories, the authors first similarly selected three broad categories on which children’s stories can be written: moral values, stories with specific characters, and stories with specific settings. They then used online websites and children’s story books to come up with topics in these areas creating a list of 35 topics. This is the topic list. Next, these were then used in a simple template ‘Explain {topic} to a / an {nationality} person.’ for QA and ‘Write a story about {topic} for a / an {nationality} kid.’ in the story. The resulting prompts were input into 6 LLMs listed in 4.2 to obtain model outputs. 5 responses were generated for every output.
The entire data of the topic list is human-curated. The model outputs are LLMs generated. Some characteristics of the model outputs are evaluated in the paper.
The cost of hand-curating topic lists was about 10 researcher hours. For getting model outputs, A6000 GPUs were used for hosting the LLM to run inference for obtaining model outputs. The total inference cost was about 45 GPU hours. Model outputs from GPT 3.5 cost about 125 USD.
We did not sample.
The data was hand-curated by the author and the author queried LLMs for model outputs. No additional personnel was involved.
No
NA
NA
NA
NA
NA
Any other comments?
NA
D.4 Preprocessing / Labelling / Cleaning
No
No cleaning was performed.
NA
Any other comments?
NA
D.5 Uses
Yes. The data, paper, and code will be open-sourced after peer review.
What (other) tasks could the dataset be used for?
No. We do note though that the model outputs are generated content from LLMs and might contain toxic, offensive, and stereotypical texts against marginalized communities. We advise discretion on the part of users who choose to further utilise this data for analysis.
The topic lists should not be treated as an exhaustive list of topics to evaluate cultural competence. The model outputs should not be used as gold standard answers for particular questions or story-generation tasks.
Any other comments?
No
D.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of the dataset was created? If so, please provide a description.
We make our data and code publicly available for replicating results and use in further analysis.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)? We make our data and code publicly available for replicating results and use in further analysis.
When will the dataset be distributed?
We make our data and code publicly available for replicating results and use in further analysis.
We make our data and code publicly available for replicating results and use in further analysis.
No
No
Any other comments?
YOUR ANSWER HERE
D.7 Maintenance
Who is supporting/hosting/maintaining the dataset?
Anonymized for peer review.
Anonymized for peer review.
Is there an erratum? If so, please provide a link or other access point.
No. The authors can be contacted via email.
Will the dataset be updated (e.g., to correct labelling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
This data is unlikely to be updated.
NA
We do not intend to have multiple versions.
TBD
Any other comments?
NA